# Asymptotic Behavior of Integral Projection Models via Genealogical Quantities

### v2

@@ 4.2. The Dominant Root and Its Eigenstructure
(86) $+\sum_{\ell=1}^{n-1}(-1)^{\ell}\sum_{k=\ell}^{n-1}\psi^{(n-k)}(x,y;r_{0})\,\widehat{B}_{\,k,\,\ell}\!\Big(\psi^{(1)},\psi^{(2)},\dots,\psi^{(k)};y,y\Big)\Bigr|_{r=r_{0}}.$
@@
(88) $w_{0}(x,y)=w_{0}(y)\left(\sum_{n=1}^{\infty}\psi_{n}(x,y;r_{0})\right),\quad w_{0}(y)\neq 0.$
@@
$\psi_{m}^{*}(y,x;r_{0}):=\$$\psi^{(m)}(y,x;r_{0})$
(89) $+\sum_{\ell=1}^{m-1}(-1)^{\ell}\sum_{k=\ell}^{m-1}\widehat{B}_{\,k,\,\ell}\!\Big(\psi^{(1)},\psi^{(2)},\dots,\psi^{(k)};y,y\Big)\Bigr|_{r=r_{0}}\,\psi^{(m-k)}(y,x;r_{0}),$
@@
(90) $v_{0}(y,x)=v_{0}(y)\left(\sum_{m=1}^{\infty}\psi_{m}^{*}(y,x;r_{0})\right),\quad v_{0}(y)\neq 0.$
@@ 4.3. Asymptotics and Demographic Interpretation
$w_{0}(a,x,y):=\exp\{-r_{0}a\}\int_{\Omega^{d}}K(a,x\leftarrow 0,\xi)\,w_{0}(\xi,y)\,d\xi.$
Accordingly, (88) expresses the stable state distribution at age zero: $w_{0}(0,x,y)=w_{0}(x,y)$.
In the simple IPM (39), the direct contribution is a discrete sum over time steps. In the multistate McKendrick equation with continuous age, the kernel $\psi(x,y;r_{0})$ represents the total lifetime reproductive contribution of an individual starting from state $y$ to offspring with initial state $x$. The functions $\psi_{n}$ in (86), which constitute $w_{0}$, represent the contribution of each generation (index $n$).
@@
(92) $\langle v_{0},\hat{G}\left(r_{0}\right)\rangle_{x}$ $=\int_{\Omega^{d}}v_{0}(y,x)\,\hat{G}\left(x;r_{0}\right)\,dx$
$=\int_{\Omega^{d}}v_{0}(y,x)\int_{0}^{\alpha}\exp\{-r_{0}a\}\int_{a}^{\alpha}\int_{\Omega^{d}}\int_{\Omega^{d}}$
(93) $\quad\times F(x\leftarrow\xi;s)\,K(s,\xi\leftarrow s-a,\eta)\,\varphi(s-a,\eta)\,d\eta\,d\xi\,ds\,da\,dx$
$=\int_{\Omega^{d}}v_{0}(y,x)\int_{0}^{\alpha}\int_{0}^{s}\exp\{-r_{0}a\}\int_{\Omega^{d}}\int_{\Omega^{d}}$
(94) $\quad\times F(x\leftarrow\xi;s)\,K(s,\xi\leftarrow s-a,\eta)\,\varphi(s-a,\eta)\,d\eta\,d\xi\,da\,ds\,dx$
$=\int_{0}^{\alpha}\int_{\Omega^{d}}\int_{\Omega^{d}}v_{0}(y,x)\int_{0}^{s}\exp\left\{-r_{0}(s-\tau)\right\}\int_{\Omega^{d}}$
(95) $\quad\times F(x\leftarrow\xi;s)\,K(s,\xi\leftarrow\tau,\eta)\,\varphi(\tau,\eta)\,d\xi\,d\tau\,dx\,d\eta\,ds$
(96) $=\int_{0}^{\alpha}\int_{\Omega^{d}}v_{0}(\tau,y,\eta)\,\varphi(\tau,\eta)\,d\eta\,d\tau$
@@
(98) $v_{0}(a,y,x):=\int_{\Omega^{d}}v_{0}(y,\xi)\int_{a}^{\alpha}\exp\{-r_{0}(\tau-a)\}\int_{\Omega^{d}}F(\xi\leftarrow\eta;\tau)\,K(\tau,\eta\leftarrow a,x)\,d\eta\,d\tau\,d\xi.$
@@
$-\left\langle v_{0},\frac{d}{dr}\hat{\mathbf{\Psi}}(r)\big|_{r=r_{0}}w_{0}\right\rangle_{x}=-\int_{\Omega^{d}}\int_{\Omega^{d}}v_{0}\left(y,x\right)\,\left.\frac{d}{dr}\psi\left(x,\xi;r\right)\right|_{r=r_{0}}\,w_{0}\left(\xi,y\right)\,d\xi\,dx$
(99) $=-v_{0}\left(y\right)w_{0}\left(y\right)\sum_{n=1}^{\infty}\sum_{m=1}^{n}\int_{\Omega^{d}}\int_{\Omega^{d}}\psi_{n-m}^{*}\left(y,x;r_{0}\right)\,\left.\frac{d}{dr}\psi\left(x,\xi;r\right)\right|_{r=r_{0}}\,\psi_{m}\left(\xi,y;r_{0}\right)\,d\xi\,dx,$
@@
(102) $E_{n}\left(y\right):=\sum_{n=1}^{\infty}n\,\psi_{n}\left(y,y;r_{0}\right).$
@@
$\bar{g}_{\mathrm{L}}:=\frac{-1}{E_{n}\left(y\right)}\sum_{n=1}^{\infty}\sum_{m=1}^{n}\int_{\Omega^{d}}\int_{\Omega^{d}}\psi_{n-m}^{*}\left(y,x;r_{0}\right)\left.\frac{d}{dr}\psi\left(x,\xi;r\right)\right|_{r=r_{0}}\psi_{m}(\xi,y;r_{0})\,d\xi\,dx.$
Using the interpretations of (98) and (90), equation (97) yields the direct contribution $\bar{R}_{\mathrm{L}}(y)$ from the initial distribution $\varphi$ to the descendants of $y$:
$\bar{R}_{\mathrm{L}}(y):=\frac{\langle v_{0},\varphi\rangle_{a,x}}{v_{0}\left(y\right)}.$
@@
(103) $E_{n}^{c}(y)$ $:=\sum_{n=1}^{\infty}\frac{n}{\lambda_{0}^{n}}\,\psi_{n}(y,y;0),$
(104) $\bar{g}_{0}$ $:=\frac{1}{E_{n}^{c}(y)}\sum_{n=1}^{\infty}\frac{1}{\lambda_{0}^{n}}\sum_{m=1}^{n}\int_{\Omega^{d}}\int_{\Omega^{d}}\psi_{n-m}^{*}(y,x;0)\,\left.\frac{d}{dr}\psi(x,\xi;r)\right|_{r=0}\,\psi_{m}(\xi,y;0)\,d\xi\,dx,$
$\bar{R}_{0}(y)$ $:=\sum_{n=1}^{\infty}\frac{1}{\lambda_{0}^{n}}\int_{0}^{\alpha}\int_{\Omega^{d}}\int_{\Omega^{d}}\psi_{n}^{*}(y,\xi;0)$
@@
In this multistate setting, $\lambda_{0}$ corresponds to the basic (net) reproduction number (Inaba, 2017). The average life expectancy adjusted for the population growth rate is
$e_{0}\left(r_{0}\right):=\frac{\int_{0}^{\alpha}\int_{\Omega^{d}}w_{0}\left(a,x,y\right)\,dx\,da}{\int_{\Omega^{d}}w_{0}\left(0,\xi,y\right)\,d\xi}.$
As a new demographic indicator, we define the per-generation total contribution of the cohort with initial state $y$, denoted by $\Upsilon(y,r_{0})$, as follows.
(106) $\Upsilon(y,r_{0}):=\frac{\int_{\Omega^{d}}w_{0}\left(0,\xi,y\right)\,d\xi}{w_{0}\left(y\right)E_{n}\left(y\right)}.$
@@
$E_{n}\left(y\right)=E_{n}^{c}\left(y\right),\quad\bar{g}_{\mathrm{L}}=\bar{g}_{0},\quad\bar{R}_{\mathrm{L}}\left(y\right)=\bar{R}_{0}\left(y\right).$
Thus, at replacement level, the stationary population is characterized by the contribution to descendants with a specific initial state $y$, the generation interval, the life expectancy at birth, and the total contribution of the cohort with initial state $y$, consistent with the classical McKendrick/Leslie theory. Furthermore, the introduction of the average contributory generation number $E_{n}^{c}(y)$ together with $\Upsilon(y,0)$ provides genealogical resolution beyond earlier models.
@@ 4.4. Other Demographic Indicators and Statistical Quantities
Similarly, the type reproduction number is the direct contribution from an ancestor with the same initial condition $y$:
(107) $T_{y}=\sum_{n=1}^{\infty}\psi_{n}(y,y;0).$
@@
(108) $1=\sum_{n=1}^{\infty}\frac{1}{\lambda_{0}^{n}}\psi_{n}(y,y;0).$
Hence, if $\lambda_{0}=1$, then $T_{y}=1$ and $r_{0}=0$. Moreover, by the monotonicity of $\left(1-\frac{1}{\lambda}\hat{\mathbf{\Psi}}(0)\right)^{-1}$ in $\lambda$, it follows that $\lambda_{0}<1$ implies $T_{y}<1$ and $r_{0}<0$.
@@
(109) $\mathbb{P}_{\mathrm{L}}([0,a);y)$ $=\frac{1}{E_{n}(y)}\sum_{n=1}^{\infty}\sum_{m=1}^{n}\int_{0}^{a}\exp\{-r_{0}\tau\}\int_{\Omega^{d}}\int_{\Omega^{d}}\psi^{*}_{\,n-m}(y,x;r_{0})$
$\qquad\times\int_{\Omega^{d}}F(x\leftarrow\eta;\tau)\,K(\tau,\eta\leftarrow 0,\xi)\,\psi_{m}(\xi,y;r_{0})\,d\eta\,d\xi\,dx\,d\tau,$
(110) $\mathbb{P}_{0}([0,a);y)$ $=\frac{1}{E_{n}^{c}(y)}\sum_{n=1}^{\infty}\frac{1}{\lambda_{0}^{\,n}}\sum_{m=1}^{n}\int_{0}^{a}\int_{\Omega^{d}}\int_{\Omega^{d}}\psi^{*}_{\,n-m}(y,x;0)$
$\qquad\times\int_{\Omega^{d}}F(x\leftarrow\eta;\tau)\,K(\tau,\eta\leftarrow 0,\xi)\,\psi_{m}(\xi,y;0)\,d\eta\,d\xi\,dx\,d\tau,$
@@
$E_{n}(y):=\sum_{n=1}^{\infty}n\,\psi_{n}(y,y;r_{0}),\qquad E_{n}^{c}(y):=\sum_{n=1}^{\infty}\frac{n}{\lambda_{0}^{\,n}}\,\psi_{n}(y,y;0).$
@@
$\mathbb{E}_{y}[\cdot]:=\int(\cdot)\,\mathbb{P}_{\mathrm{L}}(da;y),\qquad\mathbb{E}_{y}^{c}[\cdot]:=\int(\cdot)\,\mathbb{P}_{0}(da;y).$
@@
$\Theta_{\mathrm{L}}(t;y):=\ln\mathbb{E}_{y}\!\left[\exp\{ta\}\right],\qquad\Theta_{0}(t;y):=\ln\mathbb{E}_{y}^{c}\!\left[\exp\{ta\}\right],$
and the cumulants $\kappa_{k}(y):=\partial_{t}^{k}\Theta_{\mathrm{L}}(0;y)$, $\kappa_{k}^{c}:=\partial_{t}^{k}\Theta_{0}(0;y)$ $(k\geq 1)$. Then, for $r$ near $r_{0}$ and $r$ near 0, respectively,
(111) $\ln\!\Biggl(\frac{\langle v_{0},\,\hat{\mathbf{\Psi}}(r)\,w_{0}\rangle_{x}}{v(y)\,w(y)}\Biggr)$ $=\ln E_{n}(y)\;+\;\Theta_{\mathrm{L}}\!\big(-\,(r-r_{0};y)\big)$
$=\ln E_{n}(y)\;+\;\sum_{k=1}^{\infty}\frac{\kappa_{k}}{k!}\,(-1)^{k}\,(r-r_{0})^{k},$
(112) $\ln\!\Biggl(\frac{\big\langle\bar{v}_{0},\,\frac{1}{\lambda_{0}}\hat{\mathbf{\Psi}}(r)\,\bar{w}_{0}\big\rangle_{x}}{\bar{v}(y)\,\bar{w}(y)}\Biggr)$ $=\ln E_{n}^{c}(y)\;+\;\Theta_{0}\!\big(-\,r;y\big)$
$=\ln E_{n}^{c}(y)\;+\;\sum_{k=1}^{\infty}\frac{\kappa_{k}^{c}}{k!}\,(-1)^{k}\,r^{k}.$
@@
(113) $\ln\!\Biggl(\frac{\langle v_{0},\,\hat{\mathbf{\Psi}}(r)\,w_{0}\rangle_{x}}{v(y)\,w(y)}\Biggr)=\ln E_{n}(y)\;-\;\bar{g}_{\mathrm{L}}\,(r-r_{0})\;+\;\frac{1}{2}\,\sigma_{\mathrm{L}}^{2}\,(r-r_{0})^{2}\;+\;o\!\big((r-r_{0})^{2}\big),$
(114) $\ln\!\Biggl(\frac{\big\langle\bar{v}_{0},\,\frac{1}{\lambda_{0}}\hat{\mathbf{\Psi}}(r)\,\bar{w}_{0}\big\rangle_{x}}{\bar{v}(y)\,\bar{w}(y)}\Biggr)=\ln E_{n}^{c}(y)\;-\;\bar{g}_{0}\,r\;+\;\frac{1}{2}\,\sigma_{0}^{2}\,r^{2}\;+\;o\!\big(r^{2}\big),$
@@
$\bar{g}_{\mathrm{L}}$ $:=\kappa_{1}=\mathbb{E}[a],\quad\sigma_{\mathrm{L}}^{2}:=\kappa_{2}=\mathbb{V}_{y}(a),$

### v3

@@ 4.2. The Dominant Root and Its Eigenstructure
(82) $+\sum_{\ell=1}^{n-1}(-1)^{\ell}\sum_{k=\ell}^{n-1}\psi^{(n-k)}(x,y;r_{0})\,\widehat{B}_{\,k,\,\ell}\!\Big(\psi^{(1)},\psi^{(2)},\dots,\psi^{(k)};y,y\Big)\Bigr|_{r=r_{0}},$
$\psi_{1}(x,y;r_{0})$ $:=\psi(x,y;r_{0}).$
@@
Since $\rho(\hat{\mathbf{\Psi}}(r_{0}))=1$, we impose an assumption analogous to Assumption 3.1, namely:
Assumption 4.5.
There exists $y_{0}\in\Omega^{d}$ such that
$\psi(y_{0},y_{0};r_{0})>\|\hat{\mathbf{\Psi}}(r_{0})\|_{\mathrm{op}}-1.$
This condition guarantees that the $\psi_{n}$-series generated by $\psi$ remains convergent at the critical value $\lambda=\rho(\hat{\mathbf{\Psi}}(r_{0}))=1$, even on the diagonal $(y_{0},y_{0})$.
Remark 4.6 (Interpretation in age-structured models).
In age-structured models, Assumption 3.1 has a somewhat different interpretation from the simple IPM viewpoint based on one-step state transitions (e.g., slow growth or limited movement between patches). Here the diagonal point $y_{0}$ represents both the initial state of the parent and the initial state of the offspring. In other words, Assumption 3.1 amounts to requiring that, with non-negligible weight, the offspring’s initial state is genetically determined by (or closely resembles) the parent’s initial state. This provides a biologically meaningful condition in age-structured settings.
@@
(84) $w_{0}(x,y_{0})=w_{0}(y_{0})\left(\sum_{n=1}^{\infty}\psi_{n}(x,y_{0};r_{0})\right),\quad w_{0}(y)\neq 0.$
@@
(85) $\psi_{m}^{*}(y,x;r_{0}):=\$$\psi^{(m)}(y_{0},x;r_{0})$
$+\sum_{\ell=1}^{m-1}(-1)^{\ell}\sum_{k=\ell}^{m-1}\widehat{B}_{\,k,\,\ell}\!\Big(\psi^{(1)},\psi^{(2)},\dots,\psi^{(k)};y_{0},y_{0}\Big)\Bigr|_{r=r_{0}}\,$
$\times\psi^{(m-k)}(y_{0},x;r_{0}),$
@@
(86) $v_{0}(y_{0},x)=v_{0}(y)\left(\sum_{m=1}^{\infty}\psi_{m}^{*}(y_{0},x;r_{0})\right),\quad v_{0}(y)\neq 0.$
@@ 4.3. Asymptotics and Demographic Interpretation
$w_{0}(a,x,y_{0}):=\exp\{-r_{0}a\}\int_{\Omega^{d}}K(a,x\leftarrow 0,\xi)\,w_{0}(\xi,y_{0})\,d\xi.$
Accordingly, (84) expresses the stable state distribution at age zero: $w_{0}(0,x,y_{0})=w_{0}(x,y_{0})$.
In the simple IPM (38), the direct contribution is a discrete sum over time steps. In the multistate McKendrick equation with continuous age, the kernel $\psi(x,y;r_{0})$ represents the total lifetime reproductive contribution of an individual starting from state $y$ to offspring with initial state $x$. The functions $\psi_{n}$ in (4.2), which constitute $w_{0}$, represent the contribution of each generation (index $n$).
@@
$\langle v_{0},\hat{G}\left(r_{0}\right)\rangle_{x}$ $=\int_{\Omega^{d}}v_{0}(y_{0},x)\,\hat{G}\left(x;r_{0}\right)\,dx$
$=\int_{\Omega^{d}}v_{0}(y_{0},x)\int_{0}^{\alpha}\exp\{-r_{0}a\}\int_{a}^{\alpha}\int_{\Omega^{d}}\int_{\Omega^{d}}$
$\quad\times F(x\leftarrow\xi;s)\,K(s,\xi\leftarrow s-a,\eta)\,\varphi(s-a,\eta)\,d\eta\,d\xi\,ds\,da\,dx$
$=\int_{\Omega^{d}}v_{0}(y_{0},x)\int_{0}^{\alpha}\int_{0}^{s}\exp\{-r_{0}a\}\int_{\Omega^{d}}\int_{\Omega^{d}}$
$\quad\times F(x\leftarrow\xi;s)\,K(s,\xi\leftarrow s-a,\eta)\,\varphi(s-a,\eta)\,d\eta\,d\xi\,da\,ds\,dx$
$=\int_{0}^{\alpha}\int_{\Omega^{d}}\int_{\Omega^{d}}v_{0}(y_{0},x)\int_{0}^{s}\exp\left\{-r_{0}(s-\tau)\right\}\int_{\Omega^{d}}$
$\quad\times F(x\leftarrow\xi;s)\,K(s,\xi\leftarrow\tau,\eta)\,\varphi(\tau,\eta)\,d\xi\,d\tau\,dx\,d\eta\,ds$
$=\int_{0}^{\alpha}\int_{\Omega^{d}}v_{0}(\tau,y_{0},\eta)\,\varphi(\tau,\eta)\,d\eta\,d\tau$
@@
(89) $\begin{split}v_{0}(a,y_{0},x):=&\int_{\Omega^{d}}v_{0}(y_{0},\xi)\int_{a}^{\alpha}\exp\{-r_{0}(\tau-a)\}\\ &\times\int_{\Omega^{d}}F(\xi\leftarrow\eta;\tau)\,K(\tau,\eta\leftarrow a,x)\,d\eta\,d\tau\,d\xi\end{split}.$
@@
$-\left\langle v_{0},\frac{d}{dr}\hat{\mathbf{\Psi}}(r)\big|_{r=r_{0}}w_{0}\right\rangle_{x}=-\int_{\Omega^{d}}\int_{\Omega^{d}}v_{0}\left(y_{0},x\right)\,\left.\frac{d}{dr}\psi\left(x,\xi;r\right)\right|_{r=r_{0}}\,w_{0}\left(\xi,y_{0}\right)\,d\xi\,dx$
(90) $=-v_{0}\left(y\right)w_{0}\left(y_{0}\right)\sum_{n=1}^{\infty}\sum_{m=1}^{n-1}\int_{\Omega^{d}}\int_{\Omega^{d}}\psi_{n-m}^{*}\left(y_{0},x;r_{0}\right)\,$
$\quad\left.\times\frac{d}{dr}\psi\left(x,\xi;r\right)\right|_{r=r_{0}}\,\psi_{m}\left(\xi,y_{0};r_{0}\right)\,d\xi\,dx,$
@@
(93) $E_{n}\left(y_{0}\right):=\sum_{n=1}^{\infty}n\,\psi_{n}\left(y_{0},y_{0};r_{0}\right).$
@@
$\bar{g}_{\mathrm{L}}:=\frac{-1}{E_{n}\left(y_{0}\right)}\sum_{n=1}^{\infty}\sum_{m=1}^{n-1}\int_{\Omega^{d}}\int_{\Omega^{d}}\psi_{n-m}^{*}\left(y_{0},x;r_{0}\right)\left.\frac{d}{dr}\psi\left(x,\xi;r\right)\right|_{r=r_{0}}\psi_{m}(\xi,y_{0};r_{0})\,d\xi\,dx.$
Using the interpretations of (89) and (86), equation (88) yields the direct contribution $\bar{R}_{\mathrm{L}}(y_{0})$ from the initial distribution $\varphi$ to the descendants of $y_{0}$:
$\bar{R}_{\mathrm{L}}(y_{0}):=\frac{\langle v_{0},\varphi\rangle_{a,x}}{v_{0}\left(y_{0}\right)}.$
@@
(94) $E_{n}^{c}(y)$ $:=\sum_{n=1}^{\infty}\frac{n}{\lambda_{0}^{n}}\,\psi_{n}(y_{0},y_{0};0),$
(95) $\bar{g}_{0}$ $:=\frac{1}{E_{n}^{c}(y)}\sum_{n=1}^{\infty}\frac{1}{\lambda_{0}^{n}}\sum_{m=1}^{n}\int_{\Omega^{d}}\int_{\Omega^{d}}\psi_{n-m}^{*}(y_{0},x;0)\,$
$\qquad\left.\times\frac{d}{dr}\psi(x,\xi;r)\right|_{r=0}\,\psi_{m}(\xi,y_{0};0)\,d\xi\,dx,$
$\bar{R}_{0}(y)$ $:=\sum_{n=1}^{\infty}\frac{1}{\lambda_{0}^{n}}\int_{0}^{\alpha}\int_{\Omega^{d}}\int_{\Omega^{d}}\psi_{n}^{*}(y_{0},\xi;0)$
@@
In this multistate setting, $\lambda_{0}$ corresponds to the basic (net) reproduction number Inaba (2017). The average life expectancy adjusted for the population growth rate is
$e_{0}\left(r_{0}\right):=\frac{\int_{0}^{\alpha}\int_{\Omega^{d}}w_{0}\left(a,x,y_{0}\right)\,dx\,da}{\int_{\Omega^{d}}w_{0}\left(0,\xi,y_{0}\right)\,d\xi}.$
As a new demographic indicator, we define the per-generation total contribution of the cohort with initial state $y_{0}$, denoted by $\Upsilon(y_{0},r_{0})$, as follows.
(97) $\Upsilon(y_{0},r_{0}):=\frac{\int_{\Omega^{d}}w_{0}\left(0,\xi,y_{0}\right)\,d\xi}{w_{0}\left(y_{0}\right)E_{n}\left(y_{0}\right)}.$
@@
$E_{n}\left(y_{0}\right)=E_{n}^{c}\left(y_{0}\right),\quad\bar{g}_{\mathrm{L}}=\bar{g}_{0},\quad\bar{R}_{\mathrm{L}}\left(y_{0}\right)=\bar{R}_{0}\left(y_{0}\right).$
Thus, at replacement level, the stationary population is characterized by the contribution to descendants with a specific initial state $y_{0}$, the generation interval, the life expectancy at birth, and the total contribution of the cohort with initial state $y_{0}$, consistent with the classical McKendrick/Leslie theory. Furthermore, the introduction of the average contributory generation number $E_{n}^{c}(y_{0})$ together with $\Upsilon(y_{0},0)$ provides genealogical resolution beyond earlier models.
@@ 4.4. Other Demographic Indicators and Statistical Quantities
Similarly, the type reproduction number is the direct contribution from an ancestor with the same initial condition $y_{0}$:
(98) $T_{y_{0}}=\sum_{n=1}^{\infty}\psi_{n}(y_{0},y_{0};0).$
@@
(99) $1=\sum_{n=1}^{\infty}\frac{1}{\lambda_{0}^{n}}\psi_{n}(y_{0},y_{0};0).$
Hence, if $\lambda_{0}=1$, then $T_{y_{0}}=1$ and $r_{0}=0$. Moreover, by the monotonicity of $\left(1-\frac{1}{\lambda}\hat{\mathbf{\Psi}}(0)\right)^{-1}$ in $\lambda$, it follows that $\lambda_{0}<1$ implies $T_{y_{0}}<1$ and $r_{0}<0$.
@@
(100) $\mathbb{P}_{\mathrm{L}}([0,a);y_{0})$ $=\frac{1}{E_{n}(y_{0})}\sum_{n=1}^{\infty}\sum_{m=1}^{n}\int_{0}^{a}\exp\{-r_{0}\tau\}\int_{\Omega^{d}}\int_{\Omega^{d}}\psi^{*}_{\,n-m}(y_{0},x;r_{0})$
$\qquad\times\int_{\Omega^{d}}F(x\leftarrow\eta;\tau)\,K(\tau,\eta\leftarrow 0,\xi)\,\psi_{m}(\xi,y_{0};r_{0})\,d\eta\,d\xi\,dx\,d\tau,$
(101) $\mathbb{P}_{0}([0,a);y_{0})$ $=\frac{1}{E_{n}^{c}(y_{0})}\sum_{n=1}^{\infty}\frac{1}{\lambda_{0}^{\,n}}\sum_{m=1}^{n}\int_{0}^{a}\int_{\Omega^{d}}\int_{\Omega^{d}}\psi^{*}_{\,n-m}(y_{0},x;0)$
$\qquad\times\int_{\Omega^{d}}F(x\leftarrow\eta;\tau)\,K(\tau,\eta\leftarrow 0,\xi)\,\psi_{m}(\xi,y_{0};0)\,d\eta\,d\xi\,dx\,d\tau,$
@@
$E_{n}(y_{0}):=\sum_{n=1}^{\infty}n\,\psi_{n}(y_{0},y_{0};r_{0}),\qquad E_{n}^{c}(y_{0}):=\sum_{n=1}^{\infty}\frac{n}{\lambda_{0}^{\,n}}\,\psi_{n}(y_{0},y_{0};0).$
@@
$\mathbb{E}_{y_{0}}[\cdot]:=\int(\cdot)\,\mathbb{P}_{\mathrm{L}}(da;y_{0}),\qquad\mathbb{E}_{y_{0}}^{c}[\cdot]:=\int(\cdot)\,\mathbb{P}_{0}(da;y_{0}).$
@@
$\Theta_{\mathrm{L}}(t;y_{0}):=\ln\mathbb{E}_{y_{0}}\!\left[\exp\{ta\}\right],\qquad\Theta_{0}(t;y_{0}):=\ln\mathbb{E}_{y_{0}}^{c}\!\left[\exp\{ta\}\right],$
and the cumulants $\kappa_{k}(y_{0}):=\partial_{t}^{k}\Theta_{\mathrm{L}}(0;y_{0})$, $\kappa_{k}^{c}:=\partial_{t}^{k}\Theta_{0}(0;y_{0})$ $(k\geq 1)$. Then, for $r$ near $r_{0}$ and $r$ near 0, respectively,
(102) $\ln\!\Biggl(\frac{\langle v_{0},\,\hat{\mathbf{\Psi}}(r)\,w_{0}\rangle_{x}}{v(y_{0})\,w(y_{0})}\Biggr)$ $=\ln E_{n}(y_{0})\;+\;\Theta_{\mathrm{L}}\!\big(-\,(r-r_{0};y_{0})\big)$
$=\ln E_{n}(y_{0})\;+\;\sum_{k=1}^{\infty}\frac{\kappa_{k}}{k!}\,(-1)^{k}\,(r-r_{0})^{k},$
(103) $\ln\!\Biggl(\frac{\big\langle\bar{v}_{0},\,\frac{1}{\lambda_{0}}\hat{\mathbf{\Psi}}(r)\,\bar{w}_{0}\big\rangle_{x}}{\bar{v}(y_{0})\,\bar{w}(y_{0})}\Biggr)$ $=\ln E_{n}^{c}(y_{0})\;+\;\Theta_{0}\!\big(-\,r;y_{0}\big)$
$=\ln E_{n}^{c}(y_{0})\;+\;\sum_{k=1}^{\infty}\frac{\kappa_{k}^{c}}{k!}\,(-1)^{k}\,r^{k}.$
@@
(104) $\ln\!\Biggl(\frac{\langle v_{0},\,\hat{\mathbf{\Psi}}(r)\,w_{0}\rangle_{x}}{v(y_{0})\,w(y_{0})}\Biggr)=\ln E_{n}(y_{0})\;-\;\bar{g}_{\mathrm{L}}\,(r-r_{0})\;+\;\frac{1}{2}\,\sigma_{\mathrm{L}}^{2}\,(r-r_{0})^{2}\;+\;o\!\big((r-r_{0})^{2}\big),$
(105) $\ln\!\Biggl(\frac{\big\langle\bar{v}_{0},\,\frac{1}{\lambda_{0}}\hat{\mathbf{\Psi}}(r)\,\bar{w}_{0}\big\rangle_{x}}{\bar{v}(y_{0})\,\bar{w}(y_{0})}\Biggr)=\ln E_{n}^{c}(y_{0})\;-\;\bar{g}_{0}\,r\;+\;\frac{1}{2}\,\sigma_{0}^{2}\,r^{2}\;+\;o\!\big(r^{2}\big),$
@@
$\bar{g}_{\mathrm{L}}$ $:=\kappa_{1}=\mathbb{E}[a],\quad\sigma_{\mathrm{L}}^{2}:=\kappa_{2}=\mathbb{V}_{y_{0}}(a),$
@@
4.5. Consistency of the Reference-Point Normalization with the Euler–Lotka Equation
In the main text we determine the eigenvalue by normalizing the eigenfunctions so that their values at the reference point equal $1$. For the classical one-state McKendrick–von Foerster model, taking the reference point at age 0 shows that this normalization reproduces the Euler–Lotka equation.
Let $\ell(a)$ be the survival function and $\beta(a)$ the fertility rate, and write $\lambda=e^{r}>0$. In the classical theory, the stable age density has the form
$w(a)=c_{w}\,e^{-ra}\ell(a),\qquad a\geq 0,$
so that its value at age 0 is $w(0)=c_{w}$. Hence the newborn production (the boundary functional evaluated at the stable profile) is
(106) $w(0)=\int_{0}^{\infty}\beta(a)\,w(a)\,da=c_{w}\int_{0}^{\infty}e^{-ra}\beta(a)\ell(a)\,da.$
Therefore, imposing the reference-point normalization $w(0)=1$ is equivalent to
$\int_{0}^{\infty}e^{-ra}\beta(a)\ell(a)\,da=1,$
which is exactly the Euler–Lotka equation.
Likewise, under the same reference-point viewpoint, the reproductive value at age 0 is a constant multiple of the Euler–Lotka functional:
$v(0)=c_{v}\int_{0}^{\infty}e^{-ra}\beta(a)\ell(a)\,da,$
for a constant $c_{v}>0$ depending only on the chosen normalization. Recalling the scalar identity (14),
$c(\lambda)D(\lambda)=1-\frac{w(0)}{c_{w}},$
we see that, in the classical McKendrick theory, choosing age 0 as the reference point amounts precisely to the reference-point eigenstructure: the eigenvalue is recovered by fixing the eigenfunction value at the reference point.